OpenStudy (anonymous):

What is the slope of a line passing through the points A (-3,3) and B (7,-3)?

OpenStudy (anonymous):

you want to use y2-y1/x2-x1 so you would use -3-3/7+3 so you would get -6/10 which can simplify to -3/5

OpenStudy (anonymous):

Le=-3/5 e:y=-3/5x+21/5